OpenStudy (anonymous):

A rectangular garden has a perimeter of 172 feet. The length of the garden is 24 feet more than the width. What are the dimensions of the garden? (Recall the formula P = 2L + 2W.) a. 62 ft by 86 ft c. 55 ft by 79 ft b. 31 ft by 55 ft d. 74 ft by 98 ft

OpenStudy (anonymous):

this is math models btw guys. any help is much appreciated :)

OpenStudy (jdoe0001):

 \(\bf perimeter = 2l + 2w\qquad width = w\qquad length = w+24\\ \quad \\ \quad \\ 172 = 2(w)+2(w+24) \) solve for "w" to find the width, once you find the width, subtract 24, and that's the length

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw